What is the future value of 24 periodic payments of $5,070 each made at the beginning of each period and compounded at 8%? (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

What is the present value of $2,810 to be received at the beginning of each of 29 periods, discounted at 5% compound interest? (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

What is the future value of 16 deposits of $2,570 each made at the beginning of each period and compounded at 10%? (Future value as of the end of the 16th period.) (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)
Answer #1

1.Future value of annuity due=(1+rate)*Annuity[(1+rate)^time period-1]/rate

=1.08*5070[(1.08)^24-1]/0.08

=5070*72.10594

=$365,577(Approx).

2.Present value of annuity due=(1+rate)*Annuity[1-(1+interest rate)^-time period]/rate

=1.05*2810[1-(1.05)^-29]/0.05

=2810*15.89813

=$44674(Approx).

3.Future value of annuity due=(1+rate)*Annuity[(1+rate)^time period-1]/rate

=1.1*2570[(1.1)^16-1]/0.1

=2570*39.54470

=$101,630(Approx).